Question 14

1 Point

In what pH range could hydrogen peroxide (H2O2) be used as a buffer component? (Ka = 2.24 x 10-12)

Select the correct answer

  1. pH 1.35 - 3.35

  2. pH 10.6 - 12.6

  3. pH 7.15 - pH 9.15

  4. pH 2.24 - pH 4.24
Answer #1

ica= 2.24x10 12 pka = - Log (kal - = -log (2.24 X10 11.65 11.6 pkaa 11.6 pH range = pka - I to pka + 1 - = 11. 6-T 10.6 to to